OpenStudy (anonymous):

Which graph represents the function y = ½x + 4½ − 3?

OpenStudy (solomonzelman):

Assuming that what you wrote is the CORRECT problem, I would simplify the "4½−3" which would give me, 1½ . Then I would re-write it. \(\Large\color{black}{ \bf y=\frac{1}{2}x +1\frac{1}{2} }\)

OpenStudy (solomonzelman):

Now, you would have to plug in a couple of points for x. \(\Large\color{black}{ \bf y=\frac{1}{2}x +1\frac{1}{2} }\) plugging in 2 for x, \(\Large\color{black}{ \bf y=\frac{1}{2}(2) +1\frac{1}{2} }\) \(\Large\color{black}{ \bf y=1+1\frac{1}{2} }\) \(\Large\color{black}{ \bf y=2\frac{1}{2} }\) so when x=2, y=2, 1/2 thus ( 2, 2 1/2) can you find more coordinates similarly and graph the fnction.